Problem 36 I got wrong.. :(
L has y- (intercept (0,-3) and is parallel to the line with equation y =2/3 x + 1.

Answer by tutorcecilia(2152)

If the lines are parallel they have the same slope. So the slope of both lines is m = 2/3.
.
You will need to use the slope-intercept formula: y=mx+b
y=the y point
m=the slope
x=the corresponding x point
b=the y-intercept (where the line crosses the y-axis)
Plug in the values that are given for the new line with points (0, -3)
-3= (0)(2/3) + b
Solve for b
-3=b
Plug in the slope and y-intercept into the equation of the new line:
y=(2/3)x-3
or since we knew that the slope was 2/3 and the y-intercept was (0,-3), we could have gone straigth to the y=mx+b equation and plugged in the values:
y=(2/3)x-3